INAF LBC+MODS queue observing and AO PISCES, 2012 June 17-18

Observer: C. Arcidiacono, R. Briguglio, K. Boutsia, M. Faccini, F. Vitali
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LBC + PISCES

Summary

We opened the dome ~2:15UT
Problem with guiding stars on both LBCs. There are stars on Tech CCD but the telescope is unable to guide.
Problem on the RED channel: Trackers, AO & Focus stopped without any advice. We had to run lbckill/lbcstart

At 6:00UT we see that the fwhm on the Usp images for the Nonino program are no better than what we already have. We decide to stop and switch to Pisces.
All Pisces data have as "partner" in the header Arizona. Must be fixed. After image pisces.20120618.0279 the partner is set correctly to "INAF"
During an offset in closed loop the system failed a resume: opening the preset. The new preset failed on the autogain, we received a message demon error.
The BCU 39 appears frozen (not live the ccd39), the solution was to stop and start the wfs system. Operate=OFF, wfs_stop, wfs_start.
NOTE wfs_arbitrator stop and start was not solving the problem.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Nonino_19 EGS Bbess - rsloan 40mi seeing measured on images: blue:1.2" / red 1.1" - considered within specs - COMPLETED
Nonino_19 EGS Usp - rsloan 30mi seeing measured on images: blue:1.5" / red 1.23"
Standard SA104 Us/Bbess - rsloan 10min airmass 1.3
Standard SA109 Us/Bbess - rsloan 10min airmass 2.0
Cesaroni IRAS20126 Ks, Brg, H 10min - Ks / 50min - Brg+sky / 1hr H2+sky COMPLETED

Details

1:30 10Bias_Bino taken

02:48 Starting Flat test

03:20 slew for pointing correction to ACT0188
  • dofpia ,/first ,/x2
  • lbcrangebal

03:35 Standard SA104 (airmas 1,3)

03:42 Standard SA109 (airmas 2 )

03:50 focus Ob Nonino 19B
  • dofpia ,/x2
  • dimm seeing 1.1"
  • lbcrangebal

4:02 start Ob EGS_B_R_3

04:28 dofpia

04:33 start Ob EGS_B_R_4

04:53 Trackers, AO & Focus on RED channel are lost
  • Dave had to run lbckill - lbcstart because the interface did not respond to "turn-off"
  • the last red image was not saved

05:09 EGS_focus

05:19 EGS_USDT_R_3

05:39 dimm seeing 1.4"

05:58 At the end of the first OB with Uspec the fwhm on the image is 1.3"
  • since the data is not better than what we already have, we decide to stop and switch to Pisces.

05:59 SWITCH TO PISCES

06:27 Reconfiguration complete - pointing correction and collimation
  • seeing on GCS 0.7"

06:44 Slew to Cesaroni
  • 06:50 - loop closed
  • dimm seeing 1.1"
  • 2 dither patterns + sky

07:15 repeat dither patterns + sky
  • until now, no problems with the AO - the loop remains closed

07:30 the bcu stopped - need to resent the preset
  • dimm seeing 1.4"

07:34 resend preset - but it failed giving error
  • startAO failed - thSendMessage error (-3606) - no connection with MsgD
  • repeat - failed - call Alfio
  • off/on WFS arbitrator - but it did not work
  • symptom of this error the bcu was not live
  • power off AO-WFS - this worked (the problem was that the time-out was too short and it was giving error)

07:53 resend preset
  • dimm seeing 1.1"
  • 07:57 - loop closed

08:47 according to AOS the seeing is 1.6"
  • dimm seeing 1.5"

09:00 the loop opened because of drop in the counts
  • dimm seeing 1.9"
  • resend preset - shell ripped!!

09:11 loop closed again after sending the preset
  • AOS seeing 1.7"

10:42 loop opened after doing resume because of an offset
  • resend preset - failed
  • 10:50: we had to restart the WFS like before

10:55 send an active preset to take at least some sky while the WFS is rebooting

11:09 it did not manage to close the loop because the autogain did not work properly
  • send preset again
  • dimm seeing variable between 1.5"-2.1" - one of the reasons the loop keeps opening

11:54 end of science script - Close Dome

12:10 send darks

-- RobertoSpeziali - 18 Jun 2012
Topic revision: r6 - 22 Jun 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback